OpenStudy (anonymous):

A pellet gun is fired straight downward from the edge of a cliff that is 12 m above the ground. The pellet strikes the ground with a speed of 26 m/s. How far above the cliff edge would the pellet have gone had the gun been fired straight upward?

OpenStudy (anonymous):

Please help. I am confused.

OpenStudy (john_es):

First you need is the initial velocity. So we put the kinematic equations, \[v=v_0-gt\\ y=y_0+v_0t-\frac{1}{2}gt^2\]In this case, \[v_0=v+gt\\ y=y_0+(v+gt)t-\frac{1}{2}gt^2=y_0+vt+\frac{1}{2}gt^2\]Using the data, \[0=12-26t-4.9t^2\Rightarrow t=0.427\ s\]And, \[v_0=v+gt=-26+9.8\cdot0.427=-21.82\ m/s\] For the second part, first we calculate the time the bullet will fly, \[v=0=v_0-gt\Rightarrow t=\frac{v_0}{g}=\frac{21.82}{9.8}=2.23\ s\] As you see, now the initial velocity is positive because is fired in the positive senses of the y vertical direction. Calculate the height the bullet will reach above the cliff, \[h-h_0=v_0t-\frac{1}{2}gt^2=24.3 \ m\]